OpenStudy (anonymous):

In question, (3)^1/2y^2-4y-7(3)^1/2=0, use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.

OpenStudy (anonymous):

(3)^1/2y^2-4y-7(3)^1/2=0 I can't make sense of this. can you rewrite this

OpenStudy (anonymous):

sorry...

OpenStudy (anonymous):

(3)^1/2 y^2 - 4y - 7 (3)^1/2 = 0

OpenStudy (anonymous):

b^2 - 4ac (-4)^2 - 4(3^(1/2))( - 7 (3)^1/2)=100 which is positive , which mean two solution

OpenStudy (anonymous):

ok, thansks... the y just freaked me out..